OpenStudy (anonymous):

iodine 131 half life 8 days 20 mg iodine 131 is presnt in store, it was 48 days ago from laboratory finf quantity of iodine 131 left after this period

OpenStudy (triciaal):

one approach 1/2 life is time it takes for 1/2 the mass to "disappear" formula is (don't remember) 48 days ago means 48/ 8 = 6 times 20/2 = 10 after 8 days 10/2 = 5 16 days 5/2 = 2.5 24 days 2.5/2 = 1.25 32 days 1.25/2 = 0.625 40 days 0.625/2 = 0.3125g after 48 days

OpenStudy (anonymous):

thnkoo :)